OpenStudy (anonymous):

Solve: 1+(3/x-4)=12/(x(x-4))

OpenStudy (anonymous):

[3/(x-4)]-[12/x(x-4)]=-1 (3x-12)/x(x-4)=-1 3x-12=-x(x-4) 3x+x^2-4x-12=0 x^2-x-12=0 then or \[x=(-b-\sqrt{ b^2-4ac})/2a\] \[x=(1-\sqrt{1-4(1)(-12)})/2a\] \[x=(1-\sqrt{ 1+48}/2\] \[x=1/2-7/2\] or \[x=(-b+\sqrt{ b^2-4ac})/2a\ \[x=1/2+7/2\]